### v2

@@ Abstract
We consider the embeddability problem of a graph $G$ into a two-dimensional simplicial complex $C$: Given $G$ and $C$, decide whether $G$ admits a topological embedding into $C$. The problem is NP-hard, even in the restricted case where $C$ is homeomorphic to a surface. We prove that the problem is fixed-parameter tractable in the size of the two-dimensional complex, by providing an $O(2^{\text{poly}(c)}\cdot n^{2})$-time algorithm. Moreover, we show that several known problems reduce to this one, such as the crossing number and the planarity number problems, and, under some additional conditions, the embedding extension problem.
Our approach is to reduce to the case where $G$ has bounded branchwidth via an irrelevant vertex method, and to apply dynamic programming. We do not rely on any component of the existing linear-time algorithms for embedding graphs on a fixed surface, but only on algorithms from graph minor theory.
@@ 1 Introduction
An embedding of a graph $G$ into a host topological space $X$ is a crossing-free topological drawing of $G$ into $X$. The use and computation of graph embeddings is central in the communities of computational topology, topological graph theory, and graph drawing. A landmark result is the algorithm of Hopcroft and Tarjan [19], which allows to decide whether a given graph is planar (has an embedding into the plane) in linear time. Related results include more planarity testing algorithms [31], algorithms for embedding graphs on surfaces [28, 20] and for computing book embeddings [25], Hanani-Tutte theorems [37], and the theory of crossing numbers and planarization [7].
In this paper, we describe an algorithm for deciding the embeddability of graphs into topological spaces that are, in a sense, as general as possible: two-dimensional simplicial complexes (or 2-complexes for brevity), which are made from vertices, edges, and triangles glued together. (We remark that every graph is embeddable in $\mathbb{R}^{3}$ and thus in a 3-simplex, so considering higher-dimensional simplicial complexes is irrelevant.) In a previous article, jointly written with Mohar [13], we proved that, given a graph $G$ and a 2-complex $\mathscr{C}$, one can decide whether $G$ embeds into $\mathscr{C}$ in polynomial time for fixed $\mathscr{C}$; but the algorithm has running time $f(c)\cdot n^{O(c)}$, where $n$ and $c$ are the respective sizes of $G$ and $\mathscr{C}$. Using a very different strategy, we prove in this paper that it is actually fixed-parameter tractable (FPT) in the complexity of the input complex, by providing an algorithm that is quadratic in $n$ and exponential in a polynomial in $c$.
@@
One can solve the embeddability problem of graphs into 2-dimensional simplicial complexes in $2^{\text{poly}(c)}\cdot n^{2}$, where $c$ is the number of simplices of the input 2-complex and $n$ is the total number of vertices and edges of the input graph.
An additional motivation is that some other known problems related to graph drawing and planarity can be recast (and generalized) as deciding whether the graph embeds on a certain 2-complex. This includes the crossing number problem (whether a graph can be drawn in the plane with at most $k$ crossings) [21] and the planarity number problem (whether a graph with a given set of $U$ vertices can be embedded in the plane in such a way that $U$ is covered by at most $k$ faces) [4]. These problems are NP-hard in general, and we solve them in quadratic time if $k$ is fixed, in a unified way (and sometimes faster than previously known). We can also solve embedding extension problems, namely, embedding problems in which part of the graph is pre-embedded [28, 3]. We present these results in detail, with their proofs, in Section 10.
Every surface is homeomorphic to a 2-complex, but 2-complexes are much more general than surfaces, and tools that are suitable for studying embeddability of graphs on surfaces do not generalize. For example, the set of graphs embeddable on a given 2-complex is not closed under minor, which makes many tools for dealing with graphs on surfaces unsuitable for 2-complexes. Moreover, the complexity of some topological problems increase drastically when we consider 2-complexes instead of surfaces, e.g., deciding homeomorphism (solvable in linear time for surfaces, equivalent to graph isomorphism for 2-complexes [30]), or deciding the contractibility of curves (solvable in linear time for surfaces [15, 23, 18], undecidable for 2-complexes [6, 38]). See our previous paper with Mohar [13, Introduction] for more motivations.
Comparison with previous works on surfaces.
The embeddability problem has been largely considered in the special case where the input 2-complex is (homeomorphic to) a surface. This problem is already NP-hard [39], and the existing algorithms that are fixed-parameter tractable in the genus are notoriously complicated; we review them now.
@@
Kawarabayashi, Mohar, and Reed, in an extended abstract [20], have given a simpler linear-time algorithm for this problem, with a singly-exponential dependence in the genus, but not all details are presented, which makes the approach hard to check [22, p. 3657, footnote].
@@
General graph minor theory provides an algorithm for the same purpose. The graph minor theorem by Robertson and Seymour [35] implies that, for every fixed surface $\mathscr{S}$, there is a finite list of graphs $\cal O_{\mathscr{S}}$ such that a graph $G$ can be embedded on $\mathscr{S}$ if and only if $G$ does not contain any graph in $\cal O_{\mathscr{S}}$ as a minor. Moreover, there is an algorithm that given any surface $\mathscr{S}$ (specified by its genus and orientability) outputs the list $\cal O_{\mathscr{S}}$ [2], and there is an algorithm to decide whether a graph $M$ is a minor of another graph $G$ running, for fixed $M$, in time cubic in the size of $G$ [34][14, Theorem 6.12]. These considerations thus lead to an algorithm to decide embeddability of a graph on a surface that runs, if the input surface is fixed, in cubic time in the size of the input graph.
@@
Finally, in the same vein, Kociumaka and Ma. Pilipczuk [22] have studied the following more general problem than the embeddability problem of graphs on surfaces: Given a surface $\mathscr{S}$, a graph $G$, and an integer $k\geq 0$, is it possible to remove a set $W$ of at most $k$ vertices from $G$ so that $G-W$ is embeddable on $\mathscr{S}$? They provide an algorithm that is fixed-parameter tractable in $k$ and the genus of $\mathscr{S}$, where the dependence on the genus is unspecified. In particular, as a special case, they decide the embeddability of a graph on a surface; however, they use one of the previous algorithms [28, 20] as a subroutine. The problem that we study, the embeddability of graphs on 2-complexes, is independent from the problem studied by Kociumaka and Pilipczuk, in the sense that there is, a priori, no obvious reduction from one problem to the other. However, we will reuse some ingredients from that paper.
@@
Our algorithm, restricted to the case where we want to embed graphs on surfaces, is not as efficient as the existing algorithms mentioned above, because it runs in quadratic time for fixed genus. However, we stress that it is independent from the existing algorithms for embedding graphs on surfaces; we mostly rely on algorithms from graph minor theory (see in particular Lemma 8.2) and on graph planarity testing.
Overview and structure of the paper.
@@
We use a standard strategy in graph algorithms and parameterized complexity (see, e.g., the book by Cygan, Fomin, Kowalik, Lokshtanov, Marx, Pilipczuk, Pilipczuk, and Saurabh [14, Chapter 7]): we show by dynamic programming that the problem can be solved efficiently for graphs of bounded branchwidth, and then, using an irrelevant vertex method, we prove that one can assume without loss of generality that the input graph $G$ has branchwidth bounded by a polynomial in the size of the input 2-complex. In the context of surface-embedded graphs, this paradigm has been used in the extended abstract by Kawarabayashi, Mohar, and Reed [20] and in the article by Kociumaka and Pilipczuk [22]; our algorithm takes inspiration from the former one, for the idea of the dynamic programming algorithm, and from the latter, for some arguments in the irrelevant vertex method. However, handling 2-complexes requires significantly more effort. More precisely, Theorem 1.1 follows immediately from the following two theorems.
@@
We can solve the embeddability problem of graphs into two-dimensional simplicial complexes in time $(c+w)^{O(c+w)}\cdot n$ time, where $c$ is the number of simplices of the input 2-complex, $n$ is the total number of vertices and edges of the input graph, and $w$ is its branchwidth.
@@
Given a 2-complex $\mathscr{C}$ with $c$ simplices, and a graph $G$ with $n$ vertices and edges in total, we can, in $2^{\text{poly}(c)}\cdot n^{2}$ time, do one of the following:
@@
Moreover, if $G$ embeds into $\mathscr{C}$, then under mild conditions one can compute an embedding; see Section 9.
@@
Sections 3 to 7 contain the proof of Theorem 1.2. In Section 3, we show that we can make some simplifying assumptions on the input, and present data structures for representing 2-complexes and graphs embedded on them. If we restrict ourselves to the case where the input 2-complex is homeomorphic to a surface, we essentially consider combinatorial maps of graphs on surfaces, except that the graphs need not be cellularly embedded (such a data structure is called an extended combinatorial map [12, Section 2.2]). The case of 2-complexes is largely more involved.
In Section 4, we show that if our input graph $G$ has an embedding into our input 2-complex $\mathscr{C}$, then there exists an embedding of $G$ on $\mathscr{C}$ that is sparse with respect to a branch decomposition of $G$. This means that each subgraph of $G$ induced by the leaves of any subtree of the branch decomposition can be separated from the rest of $G$ using a graph embedded on $\mathscr{C}$, called partitioning graph, of small complexity. We find that this new structural result, even in the surface case, is interesting and can prove useful in other contexts. If the target space were a surface, we could assume that $G$ is 3-connected and has no loop or multiple edge, which would imply (still with some work) that any embedding of $G$ would be sparse, but again the fact that we consider 2-complexes requires additional work.
@@
In Section 9, we explain how to compute an embedding of $G$ into $\mathscr{C}$, if one exists, under mild conditions.
Finally, we prove the promised applications in Section 10.
@@ 2.1 Graphs and branch decompositions
  • every node has degree either one (it is a leaf) or three (it is an internal node),
  • the root is a leaf,
  • each non-root leaf is labelled with an edge of $G$, and this labelling induces a bijection,
  • the two children of each non-leaf node are ordered.
@@ 2.2 Surfaces
An embedding of a graph $G$ into a surface $\mathscr{S}$ is cellular if each face of the embedding is homeomorphic to an open disk. If $G$ is cellularly embedded on a surface with genus $g$ and $b$ boundary components, such that the embedding has $v$ vertices, $e$ edges, and $f$ faces, then Euler’s formula stipulates that $v-e+f=2-g-b$ (these quantities are referred to as the Euler characteristic of the surface).
@@
A 2-complex (or two-dimensional simplicial complex) is an abstract simplicial complex of dimension at most two: a finite set of 0-simplices called vertices, 1-simplices called edges, and 2-simplices called triangles. Each edge is a pair of vertices, and each triangle is a triple of vertices; moreover, each subset of size two in a triangle must be an edge. (We slightly depart from the standard definition in the sense that, for us, a 2-complex is a simplicial complex in which each simplex has dimension at most two; there needs not be a simplex of dimension exactly two.)

### v3

@@ Abstract
We consider the embeddability problem of a graph $G$ into a two-dimensional simplicial complex $C$: Given $G$ and $C$, decide whether $G$ admits a topological embedding into $C$. The problem is NP-hard, even in the restricted case where $C$ is homeomorphic to a surface. We prove that the problem is fixed-parameter tractable in the size of the two-dimensional complex, by providing an $O(2^{\text{poly}(c)}\cdot n^{2})$-time algorithm. If $G$ embeds into $C$, we can compute a representation of an embedding in the same amount of time. Moreover, we show that several known problems reduce to this one, such as the crossing number and the planarity number problems, and, under some conditions, the embedding extension problem.
Our approach is to reduce to the case where $G$ has bounded branchwidth via an irrelevant vertex method, and to apply dynamic programming. We do not rely on any component of the existing linear-time algorithms for embedding graphs on a fixed surface, but only on algorithms from graph minor theory. However, by combining our results with a linear-time algorithm for embedding graphs on surfaces and with a very recent result for the irrelevant vertex method, we can decide whether $G$ embeds into $C$ in $f(c)\cdot O(n)$ time, for some function $f$.
Keywords: computational topology, graph algorithm, embedding, simplicial complex, graph, surface
2020 MSC Classification: 57-08 Computational methods for problems pertaining to manifolds and cell complexes; 57M15 Relations of low-dimensional topology with graph theory; 68Q27 Parameterized complexity, tractability and kernelization; 68W05 Nonnumerical algorithms
@@ 1 Introduction
An embedding of a graph $G$ into a host topological space $X$ is a crossing-free topological drawing of $G$ into $X$. The use and computation of graph embeddings is central in the communities of computational geometry/topology, algorithms for planar and surface-embedded graphs, topological graph theory, and graph drawing. A landmark result is the algorithm of Hopcroft and Tarjan [24], which allows to decide whether a given graph is planar (has an embedding into the plane) in linear time. Related results include more planarity testing algorithms [38], algorithms for embedding graphs on surfaces [35, 25] and for computing book embeddings [32], Hanani-Tutte theorems [46], and the theory of crossing numbers and planarization [45, 8].
Our result.
In this paper, we describe an algorithm for deciding the embeddability of graphs into topological spaces that are, in a sense, as general as possible: two-dimensional simplicial complexes (or 2-complexes for brevity), which are made from vertices, edges, and triangles glued together. (We remark that every graph is embeddable in $\mathbb{R}^{3}$ and thus in a 3-simplex, so considering higher-dimensional simplicial complexes is irrelevant.) In a previous article, jointly written with Mohar [15], we proved that, given a graph $G$ and a 2-complex $\mathscr{C}$, deciding whether $G$ embeds into $\mathscr{C}$ is NP-complete and is in XP: it can be done in polynomial time for fixed $\mathscr{C}$, but the algorithm has running time $f(c)\cdot n^{O(c)}$, where $n$ and $c$ are the respective sizes of $G$ and $\mathscr{C}$. Using a very different strategy, we prove in this paper that it is actually fixed-parameter tractable (FPT) in the complexity of the input complex, by providing an algorithm that is quadratic in $n$ and exponential in a polynomial in $c$:
@@
One can solve the embeddability problem of graphs into 2-dimensional simplicial complexes in time $O(2^{\text{poly}(c)}\cdot n^{2})$, where $c$ is the number of simplices of the input 2-complex and $n$ is the total number of vertices and edges of the input graph.
(In our proof, the polynomial in $c$ has degree $18+\varepsilon$, for any $\varepsilon>0$.) We can also compute a representation of an embedding if there is one; see Section 9 for details.
We remark that with additional ingredients, namely, a very recent paper [23] and an algorithm to embed graphs on surfaces [35], we can obtain an alternative decision algorithm that runs in time $f(c)\cdot O(n)$, for some unspecified function $f$. However, we only discuss this in the conclusion, Section 11. Indeed, as we shall see, one of the benefits of our algorithm for Theorem 1.1 is that it does not rely on any of the (complicated) algorithms for embedding graphs on surfaces, and actually provides a simpler one, at the cost of a worse dependency in the input graph size $n$.
Applications.
Our embeddability problem is quite general and captures some known parameterized problems related to graph drawing and planarity, which are NP-hard in general, and we solve them in quadratic time, in a unified way.
The most relevant problem in this vein is the crossing number problem, namely, to decide whether a graph can be drawn in the plane with at most $k$ crossings. The problem is NP-complete [22]. As is already known [15, Introduction], there is an easy reduction from the crossing number problem to the embeddability problem of graphs in 2-complexes, which immediately implies a $2^{\text{poly}(k)}\cdot O(n^{2})$-time algorithm for the crossing number problem. Previously, no FPT algorithm with runtime singly-exponential in $k$ was known, although since the conference version of this paper, Lokshtanov, Panolan, Saurabh, Sharma, Xue, and Zehavi, in a very recent breakthrough paper [31], have given a $k^{O(k)}\cdot n$-time algorithm for the crossing number problem.
In a similar manner, the planarity number problem (whether a graph with a given set $U$ of vertices can be embedded in the plane in such a way that $U$ is covered by at most $k$ faces) has an easy reduction to our problem, implying that it is FPT in $k$, although the latter fact was proved earlier using very different techniques [5].
Actually, our techniques extend immediately to new, more general versions of all of these problems in which the plane is replaced with a fixed surface of genus $g$; in that case, the algorithms are FPT in $k+g$. We can also solve embedding extension problems, namely, embedding problems in which part of the graph is pre-embedded [35, 3].
All these results are detailed and proved in Section 10. We also remark that, since the conference version of this paper appeared, more applications of Theorem 1.1 have been found; as it turns out, many kinds of crossing number problems encountered in graph drawing reduce to the embeddability of graphs on 2-complexes, so that our result applies to several other well-studied crossing number variants, for which fixed-parameter tractability parameterized by the total number of crossings was not known [14]. Thus, our algorithm appears to be a versatile and powerful tool that allows to solve diverse problems in a relatively simple way.
Embedding graphs on surfaces.
Our embeddability problem also contains, as a special case, the problem of embedding a graph into a surface, since every surface of genus $g$ (a “sphere with $g$ handles or $g$ crosscaps”) is (homeomorphic to) some 2-complex with size $O(g)$. This problem is of paramount importance; indeed, many graph problems turn out to be more easily solvable for graphs embedded on low-genus surfaces than for general graphs (see, e.g., the survey [12, Section 6]), so an embedding algorithm yields faster graph algorithms for abstract graphs that happen to be embeddable on low-genus surfaces. The embedding problem is already NP-hard for surfaces [48], and the existing algorithms that are fixed-parameter tractable in the genus are notoriously complicated; we review them now.
@@
Kawarabayashi, Mohar, and Reed, in an extended abstract [25], have given a simpler linear-time algorithm for this problem, with a singly-exponential dependence in the genus, but not all details are presented, which makes the approach hard to check [27, p. 3657, footnote].
@@
General graph minor theory provides an algorithm for the same purpose. The graph minor theorem by Robertson and Seymour [42] implies that, for every fixed surface $\mathscr{S}$, there is a finite list of graphs $\cal O_{\mathscr{S}}$ such that a graph $G$ can be embedded on $\mathscr{S}$ if and only if $G$ does not contain any graph in $\cal O_{\mathscr{S}}$ as a minor. Moreover, there is an algorithm that given any surface $\mathscr{S}$ (specified by its genus and orientability) outputs the list $\cal O_{\mathscr{S}}$ [2]. Also, there is an algorithm to decide whether a fixed graph is a minor of an input graph $G$ running in $O(n^{1+o(1)})$ time, where $n$ is the size of $G$ [28] (improving over cubic-time algorithms [41][17, Theorem 6.12]). These considerations thus lead to an algorithm to decide embeddability of a graph on a surface that runs, if the input surface is fixed, in almost-linear time in the size of the input graph. However, with this approach, it is not possible to compute an embedding if one exists.
@@
Finally, in the same vein, Kociumaka and Ma. Pilipczuk [27] have studied the following problem, which is more general than the embeddability problem of graphs on surfaces: Given a surface $\mathscr{S}$, a graph $G$, and an integer $k\geq 0$, is it possible to remove a set $W$ of at most $k$ vertices from $G$ so that GWG-W is embeddable on $\mathscr{S}$? They provide an algorithm that is fixed-parameter tractable in $k$ and the genus of $\mathscr{S}$, where the dependence on the genus is unspecified. However, they use one of the previous algorithms for embedding graphs on surfaces [35, 25] as a subroutine. The problem that we study, the embeddability of graphs on 2-complexes, is independent from the problem studied by Kociumaka and Pilipczuk, in the sense that there is, a priori, no obvious reduction from one problem to the other. However, we will reuse some ideas from that paper.
@@
Our algorithm, restricted to the case where we want to embed graphs on surfaces, is not as efficient as the existing algorithms mentioned above, because it runs in quadratic time for fixed genus. However, we stress that it does not rely on the toolbox used in the previous algorithms for embedding graphs on surfaces; we essentially rely on algorithms from graph minor theory (see in particular Lemma 8.2) and on graph planarity testing.
Comparison between surfaces and 2-complexes.
Every surface is homeomorphic to a 2-complex, but 2-complexes are much more general than surfaces, and tools that are suitable for studying embeddability of graphs on surfaces do not generalize. For example, the set of graphs embeddable on a given 2-complex is not closed under minor, which makes many tools for dealing with graphs on surfaces unsuitable for 2-complexes. Moreover, the complexity of some topological problems increase drastically when we consider 2-complexes instead of surfaces, e.g., deciding homeomorphism (solvable in linear time for surfaces, equivalent to graph isomorphism for 2-complexes [37]), or deciding the contractibility of curves (solvable in linear time for surfaces [18, 29, 21], undecidable for 2-complexes [7, 47]). See our previous paper with Mohar [15, Introduction] for more details. In this sense, it is perhaps surprising to obtain an FPT algorithm for embedding graphs on 2-complexes. We also remark that the class of 2-complexes includes the pseudosurfaces considered in graph theory; see Archdeacon [4, Section 5.7] for a survey.
Overview.
@@
At a very high level, we use a standard strategy in graph algorithms and parameterized complexity (see, e.g., the book by Cygan, Fomin, Kowalik, Lokshtanov, Marx, Pilipczuk, Pilipczuk, and Saurabh [17, Chapter 7]): we show by dynamic programming that the problem can be solved efficiently for graphs of bounded branchwidth, and then, using an irrelevant vertex method, we prove that one can assume without loss of generality that the input graph $G$ has branchwidth bounded by a polynomial in the size of the input 2-complex. In the context of surface-embedded graphs, this paradigm has been used in the extended abstract by Kawarabayashi, Mohar, and Reed [25] and in the article by Kociumaka and Pilipczuk [27]; our algorithm takes inspiration from the former one, for the idea of the dynamic programming algorithm, and from the latter, for some arguments in the irrelevant vertex method. However, handling 2-complexes requires significantly more effort, new data structures, and some tailored preprocessing steps. More precisely, Theorem 1.1 follows immediately from the following two theorems:
@@
We can solve the embeddability problem of graphs into two-dimensional simplicial complexes in time $(c+w)^{O(c+w)}\cdot n$, where $c$ is the number of simplices of the input 2-complex, $n$ is the total number of vertices and edges of the input graph, and $w$ is its branchwidth.
@@
Given a 2-complex $\mathscr{C}$ with $c$ simplices, and a graph $G$ with $n$ vertices and edges in total, we can, in time $O(2^{\text{poly}(c)}\cdot n^{2})$, do one of the following:
@@
(In our proof, the polynomial in $c$ has degree $18+\varepsilon$, for any $\varepsilon>0$.)
Incidentally, Theorem 1.2 (together with the reduction presented in Section 10.1) immediately implies that the crossing number $k$ of a graph of branchwidth (or, equivalently, treewidth) $w$ can be computed in $(k+w)^{O(k+w)}\cdot n$ time, and this improves over the best previous result [26], in which the dependence on $k+w$ is triply exponential, since it relies on Courcelle’s theorem [16]. As mentioned above, in a very recent paper [31], Lokshtanov, Panolan, Saurabh, Sharma, Xue, and Zehavi provide an algorithm for computing the crossing number $k$ of arbitrary graphs in $k^{O(k)}\cdot n$ time. To obtain this landmark result, they critically rely on an algorithm for the crossing number problem that runs in $(k+w)^{O(k+w)}\cdot n$ time. While, for technical reasons in order to be reusable for the algorithm for arbitrary branchwidth, their bounded-branchwidth algorithm solves a more general problem than the crossing number problem, the present paper is the first to present such an algorithm.
Structure of the paper.
@@
Sections 3 to 7 contain the proof of Theorem 1.2. In Section 3, we show that we can make some simplifying assumptions on the input. In particular, every graph embeds in a 3-book (a 2-complex obtained by identifying three triangles along a common edge), so in the rest of the paper we assume that the input 2-complex has no 3-book. We then present data structures for representing 2-complexes and graphs embedded on them. If we restrict ourselves to the case where the input 2-complex is homeomorphic to a surface, we essentially consider combinatorial maps of graphs on surfaces, except that the graphs need not be cellularly embedded (such a data structure is called an extended combinatorial map [13, Section 2.2]). The case of 2-complexes is largely more involved.
In Section 4, we show that if our input graph $G$ has an embedding into our input 2-complex $\mathscr{C}$, then there exists an embedding of $G$ on $\mathscr{C}$ that is sparse with respect to a branch decomposition of $G$. This means that each subgraph of $G$ induced by the leaves of any subtree of the branch decomposition can be separated from the rest of $G$ using a graph embedded on $\mathscr{C}$, called partitioning graph, of small complexity. We find that this new structural result, even in the surface case, is interesting and might prove useful in other contexts. If the target space were a surface, we could assume that $G$ is 3-connected and has no loop or multiple edge, which would imply (still with some work) that any embedding of $G$ would be sparse, but again the fact that we consider 2-complexes requires additional work.
@@
In Section 9, we explain how to compute an embedding of $G$ into $\mathscr{C}$, if one exists. We prove the promised applications in Section 10. Finally, we conclude in Section 11.
@@ 2.1 Graphs and branch decompositions
  • the root has a single child,
  • every non-root node is either an internal node, in which case it has exactly two children, which are ordered, or a leaf, in which case it has no child,
  • every (non-root) leaf is labelled with an edge of $G$, and this labelling induces a bijection.
@@ 2.2 Surfaces
An embedding of a graph $G$ into a surface $\mathscr{S}$ is cellular if each face of the embedding is homeomorphic to an open disk. If $G$ is cellularly embedded on a surface with genus $g$ and $b$ boundary components, such that the embedding has $v$ vertices, $e$ edges, and $f$ faces, then Euler’s formula stipulates that ve+f=2gbv-e+f=2-g-b (these quantities are referred to as the Euler characteristic of the surface).
@@
A 2-complex (or two-dimensional simplicial complex) is an abstract simplicial complex of dimension at most two: a finite set of 0-simplices called vertices, 1-simplices called edges, and 2-simplices called triangles. Each edge is a pair of vertices, and each triangle is a triple of vertices; moreover, each subset of size two in a triangle must be an edge. The 0-, 1-, and 2-simplices form the simplices of the complex. (We slightly depart from the standard definition in the sense that, for us, a 2-complex is a simplicial complex in which each simplex has dimension at most two; there needs not be a simplex of dimension exactly two.)